OpenStudy (anonymous):

x^2-9x+8 Polynomial expression

OpenStudy (anonymous):

Yes it is.. :)

OpenStudy (anonymous):

Need help

OpenStudy (anonymous):

did u want that polynomial factored?

OpenStudy (anonymous):

yes @ByteMe

OpenStudy (anonymous):

First multiply 8 by x^2 and broke it down to the multiplication of two numbers which add up to -9x...

OpenStudy (anonymous):

-3 times -3 equal -9

OpenStudy (anonymous):

\[8x^2=-4x \qquad 2x = -6x\]\[8x^2=8x \qquad 1x = 9x\] Likewise continue till you get a sum of -9x..

OpenStudy (anonymous):

-3 *-3=+9

OpenStudy (anonymous):

oh okay but dont we have to get 8

OpenStudy (anonymous):

\[8x^2=-8x\quad-x=-9x\]

OpenStudy (anonymous):

Use that..

OpenStudy (anonymous):

think of two integers that : add up to -9 and multiply to +8

OpenStudy (anonymous):

So you get \[\large\color{green}{x^2-8x-x+8}\]

OpenStudy (anonymous):

That is a horrible way of solving it because it simply does not work when you are dealing with complicated functions. Basically, you want the points were the function is 0. So you do the following to complete the square: \[ax^2+bx+c=0\]Multiply by 4a\[4a^2x^2+4abx+4ac=0\]Sum b^2\[4a^2x^2+4abx+4ac+b^2=b^2\]Complete the square:\[(2ax+b)^2+4ac=b^2\]\[(2ax+b)^2=b^2-4ac\]\[2ax+b=\pm \sqrt{b^2-4ac}\]\[x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\]From that you get 2 values for wich x is zero, and the factorization is: (x-x1)(x-x2) where x1 and x2 must be the points were the function is 0.

OpenStudy (anonymous):

Sorry, on the first line of the last paragraph I meant f is 0 not x.